Shifter leader was sjs Shifters were RHIC cryo problems for much of shift. Started shift at an end of a store. With FV and HI, ran a number of test runs in an attempt to understand the apparent loss of efficiency for peripheral events seen in the 200 GeV CuCu data. No obvious trigger or timing problems found. One good data run during shift (about two hours): Run 13860 (5.66K trig 2). What was intended as a data run, 13858, was started with the run type incorrectly left on pedestal. This crashed the ioserver. No new detector problems seen. _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Sat Mar 19 00:04:49 2005
This archive was generated by hypermail 2.1.8 : Sat Mar 19 2005 - 00:05:03 EST